# Impact of domain reduction techniques in polynomial optimization: A computational study

### v1

@@ 3.2 FBBT enhancements based on duality
We start with the approach proposed in [Gleixner:2017]. After solving all OBBT’s problems we have, for each variable, one solution for its maximization problem and another one for its minimization problem. Suppose we have the primal-dual solution from solving the minimization problem for variable , where is the vector of multipliers associated with the constraints of the OBBT relaxation and multiplier corresponds with an additional constraint bounding the objective function to the best upper bound available when OBBT was applied. Given a valid upper bound for (PO), the following constraint is a valid cut for (PO):
where is the vector of right-hand sides of the linear relaxation used by OBBT. Similarly, we can use the maximization problem to obtain an upper bound for variable .
Now, we describe the cuts presented in [Ryoo:1996]. Consider the relaxation of (PO) at a given node and suppose that constraint is active at the solution of this node. Suppose, moreover, that this constraint has multiplier and that is a valid upper bound for (PO). Then, the following constraint is a valid cut for (PO):
If we had an equality constraint, we would use the sign of to determine the sense of the new cut’s inequality. This is also applied to the constraints of the form and (where and are the upper and lower bounds of variable ), obtaining
It is worth emphasizing that the values for the bounds and in constraints (5) and (6) are not static: they are updated as the algorithm progresses. Put differently, whenever any of these constraints are added to the FBBT at a given node, the current values for the bounds are used.
@@ 3.2.1 Numerical results
In order to assess the performance of the aforementioned approaches, we have studied different variations, depending on how aggressively we use them. We present here the best performing ones:
@@
FBBT: The OBBT-derived cuts from [Gleixner:2017] are added to FBBT at the root node and at nodes whose depth is a multiple of 50.
FBBT: The node by node cuts from [Ryoo:1996] are added to FBBT at the root node and at nodes whose depth is a multiple of 10.
FBBT: Both approaches are applied together.
In Section A in the Appendix we report some additional results which, in particular, show that applying these enhancements at all nodes is not desirable. This is because the computational overhead required to perform FBBT’s constraint propagation on the extended set of constraints is not compensated by the gains from bound reduction. FBBT and FBBT provide a good balance between the computational effort and the reduction of the size of the branch-and-bound tree.
@@
FBBT | 287 | |||||
FBBT | 287 | |||||
FBBT | 287 |
@@ 3.3 Impact of the branching point on the branch-and-bound tree
Whenever a subproblem is solved at a node of the branch-and-bound tree, important decisions have to be made. Two crucial ones are the selection of the branching variable and of the precise value of that variable on which branching is done. As mentioned in the introduction, the choice of the branching variable has been thoroughly discussed by past literature, mainly for MILP problems (see, for instance, [Linderoth1999] and [Achterberg2005]), but also in the context of nonlinear optimization and spatial branching in [Tawarmalani2004] and [Belotti2009]. When it comes to polynomial optimization and the RLT technique in particular, [Gonzalez-Rodriguez:2023] and [Ghaddar:2023] offer deep analyses on the impact of variable selection.
@@ 3.3.1 Numerical results
Given the results in Table 4, we see that OV and MP are particularly bad at Gap and Time, respectively. Thus, one may wonder whether the superior performance of the strict convex combinations comes from the fact they just avoid sticking to OV or MP in problems where they do not perform well. Put differently, would a hypothetical version of the solver capable of choosing the best performing approach between OV and MP outperform the strict convex combinations? The answer to this question is negative. Table 5 shows, for instance, that the aforementioned hypothetical configuration would be within 5% of the best performing approach in Time in, at most, 96 out of the 138 reported instances, whereas approach 0.5OV+0.5MP on its own is within 5% of the best performing one in 102 instances.
(403 instances) | 5% Gap (110) | 5% Time (138) | 5% Pace (250) | 5% Nodes (280) |
@@ 3.4 Impact of combined enhancements
As a complement to the individual analysis in the preceding subsections, regarding the performance of the different approaches to enhance domain reduction techniques, we now study the impact of incorporating all of them together. Note that, although all the enhancements deal with different aspects of domain reduction, they are essentially independent from one another, so there should be no detrimental effect on performance when combining them: i) the enhancements in Section 3.1 only affect the OBBT at the root node, ii) the enhancements in Section 3.2 just affect the FBBT at all nodes, and iii) the enhancements in Section 3.3 deal with the branching point selection, which is independent of the OBBT and FBBT functionalities. For the sake of exposition, we just report the performance of combinations of the best performing approaches for each individual enhancement: SOCP, FBBT, and 0.5OV+0.5MP. Yet, it is important to note that retrieving dual information for (nonlinear) conic problems is far from straightforward. Thus, whenever a combination uses SOCP, we replace FBBT with FBBT, the best performing approach in Section 3.2 once we disregard the approaches applying FBBT.
@@
SOCP + FBBT | 287 | |||
SOCP + 0.5OV+0.5MP | 289 | |||
FBBT + 0.5OV+0.5MP | 287 | |||
SOCP + FBBT + 0.5OV+0.5MP | 287 |
In Table 6 we can see that, as expected from the previous results, the combinations that use approach 0.5OV+0.5MP are the most competitive ones. Among them, probably FBBT + 0.5OV+0.5MP is slightly superior to the rest, with the raw 0.5OV+0.5MP performing very similarly. Interestingly, the combination SOCP + 0.5OV+0.5MP is the one that solves most instances, despite not being so competitive in Time.
@@ 4 Improving performance with machine learning
In this section we take the enhancements of Section 3 one step further. The goal is to study the potential for additional performance improvements by using machine learning techniques to predict the best combination of enhancements for a given instance. To this end, we also analyze the improvement on performance we would observe if we had an oracle signalling the best combination for each instance and, then, we use it as a reference to assess the quality of the learning.
@@ 4.1 Machine learning framework
In [Ghaddar:2023], the authors present a machine learning framework designed to “optimally” choose between different configurations of a solver when confronting a new instance, given its underlying features. In order to do so, the machine learning model is first trained on a predefined set of instances, using both their features and the performance of the different configurations on each of them. For our analysis we follow the best performing approach among those reported in [Ghaddar:2023], choosing Pace as the measure of performance (KPI) and using quantile regression forests for the learning [Meinshausen:2006]. One of the advantages of random forests, as well as other ensemble methods that use bagging, is that we the complete data set can be used to evaluate the model using the out-of-bag predictions, without randomly splitting the data into training and test sets.222As explained in [Ghaddar:2023], given a data set of size , “a large number of bootstrap samples of size n are obtained from the data set by sampling with replacement. Each bootstrap sample leaves out, on average, about 37% of the observations (the left-out observations constitute the so-called out-of-bag sample). Then, individual trees are grown using these bootstrap samples. For each observation in the data set, the corresponding out-of-bag prediction is obtained by taking into account only those trees fitted on bootstrap samples that leave out that particular observation”. The learning is carried out on the same set of instances of the preceding section, using the statistical language R [R-Core-Team:2023] with the library ranger [Wright:2017].
@@
It is important to note that, not only we follow the approach in [Ghaddar:2023], but we also use the same set of features and the same parameters for quantile regression, so there is no tailor-made adaptation to the current setting. Thus, our results can be seen as a lower bound on what might be achieved by considering additional features that capture relevant aspects for domain reduction.
In the results’ tables in this section we report the performance of a special approach: Optimal, which represents the performance obtained when we choose, for each instance, the best configuration according to pace, i.e., it’s an ideal version representing what we would get if we had an oracle signalling the best approach for each instance. If, for a given instance, there are two options with the same pace, we choose the one with the smallest optimality gap.
@@ 4.2 Impact of machine learning on the individual enhancements
Table 7 summarizes the results of our learning methodology when applied individually to each one of the three different enhancements discussed in Section 3. More precisely, for each enhancement, the goal is to predict which one of the considered approaches will have a better performance on a given instance. To simplify the tables of results, among the original approaches we just represent the best performing one for the given enhancement, Best. “out-of-bag Q-RF” is the result of choosing, for each instance, the approach selected by the machine learning framework. Finally, “Improvement after learning” is the percentage of relative improvement from Best to out-of-bag Q-RF and “Optimal Improvement” is the same but comparing Best to Optimal.
@@
Optimal improvement |
The results in Table 7 are somewhat mixed. For the conic bound tightening from Section 3.1, we see modest improvements in Gap and Time, but relatively far from the ones delivered by Optimal. Learning seems to perform particularly well for the FBBT enhancements from Section 3.2, with out-of-bag Q-RF significantly outperforming Best and getting around half-way of the improvements Optimal would achieve. Finally, regarding the selection of the branching point from Section 3.3, the learning does not seem to obtain much since, although it improves almost by 2% in Time with respect to Best, its performance deteriorates in Gap and Pace by 0.4% and 1.19%, respectively. It seems as if the inherent randomness behind the impact of the branching point on the generated branch-and-bound trees is hard to learn upon.
@@ 4.3 Impact of machine learning on the combined enhancements
Despite the mixed results of the previous section, one would expect that, when adding all of the enhancements and its combinations together, there is more potential for learning, given the additional richness and diversity of the underlying approaches/configurations. For the sake of exposition, learning is not carried out on all approaches and combinations, but just on those reported in Section 3.4.
@@
The results in Table 8 show that machine learning performs fairly well. The learning approach outperforms Best (FBBT + 0.5OV+0.5MP) according to all the metrics: it solves three more instances and obtains improvements of around 8% and 3% in Time and Gap. As expected, having a richer set of configurations to choose from also helps Optimal to obtain larger improvements.
@@
Figure 1 represents, side by side, how often each of the eight configurations is selected by out-of-bag Q-RF and by Optimal. We can see that out-of-bag Q-RF mimics quite well the behavior of Optimal. Importantly, Figure 1 also shows that a configuration with SOCP is chosen in around one third of the instances, demonstrating the potential of using conic-based relaxations in the OBBT stages of global optimization solvers.
@@
Despite the apparently nice behavior of out-of-bag Q-RF displayed in Figure 1, it might even be the case that it never chose the best configuration and just happened to use the different configurations with the right frequencies. Figure 2 presents detailed information regarding the performance of the different configurations which, in particular, allows to assess how often each of them is the best one. The bar chart represents, for each configuration, the percentage of instances in which that configuration was the best one, the second best, and so on, for the eight ranking positions it might occupy depending on its performance relative to the rest.333Note that we have 8 configurations plus out-of-bag Q-RF, and we just consider ranking positions from 1 to 8. This is because out-of-bag Q-RF always coincides with one of the eight configurations on which the learning is performed. Moreover, the numbers inside the bars indicate how close that configuration is, on average, to the best one for the instances in which it occupies the corresponding ranking position. More specifically, these numbers go from zero to one and are computed by dividing the best (smallest) Pace among the different configurations by the Pace of the current one, and then taking the average over the instances in each ranking position. Thus, Figure 2 allows to assess not only how often each approach is in each ranking position, but also how rapidly its performance deteriorates as it goes from the top positions to the bottom ones.
@@
Figure 2 shows that out-of-bag Q-RF comes on top more frequently than any other configuration, with 0.5OV+0.5MP close behind. Yet, the main difference between them comes from the fact that the relative performance of out-of-bag Q-RF deteriorates more slowly as it moves from being the best configuration to the second or third best, for example. In Figure 2 we can also see that the configurations that use SOCP are the ones in which performance tends to deteriorate more quickly, which is consistent with the discussion in Section 3.1 about their “riskiness”. In this respect, since out-of-bag Q-RF uses SOCP in nearly one third of the instances, the fact that it performs so well even when it does not choose the top performing configuration suggests that it is successful at learning to avoid using SOCP in instances in which it does not perform well. Yet, the difference in performance between out-of-bag Q-RF and Optimal also suggests that there is room for further improvement. A natural direction would be to enhance the set of features of the instances, with the goal of capturing additional aspects that might be relevant for domain reduction.

### v2

@@ 3.2 FBBT enhancements based on duality
We start by outlining the approach proposed in [Gleixner:2017]. Suppose that is a valid upper bound for (PO), available at the moment of applying OBBT. We can tighten the OBBT relaxations by adding constraint . Then, after solving all OBBT’s problems, we have, for each variable, one solution for its maximization problem and another one for its minimization problem. Suppose we have the primal-dual solution from solving the minimization problem for variable , where is the vector of multipliers associated with the constraints of the OBBT relaxation and multiplier corresponds with constraint . Finally, let be the vector of reduced costs associated with . Theorem 1 in [Gleixner:2017] establishes that the following constraint is a valid cut for (PO):
where comes from the right-hand sides of the linear relaxation used by OBBT (including the coefficients and the right-hand sides of the bound-factor constraints). Similarly, we can use the maximization OBBT problem to obtain cut . Interestingly, if an improved solution of (PO) is found, Theorem 1 in [Gleixner:2017] also ensures that the above cuts can be tightened by updating the value of .
Now, we describe the cuts presented in [Ryoo:1996]. Consider the relaxation of (PO) at a given node. Let be its optimal value and suppose that a certain constraint is active at the solution of this node.444For instance, for the original constraints of (PO), they would be written as , where represents the linearization of . Suppose, moreover, that this constraint has associated the dual multiplier and that is a valid upper bound for (PO). Then, the following constraint is a valid cut for (PO):
The above cuts also apply to constraints of the form and (where and are the upper and lower bounds of variable at the given node), obtaining
If we had an equality constraint , we would use the sign of to determine the sense of the new cut’s inequality. Again, is worth emphasizing that the values for the bounds and in the above cuts are not static: they are updated as the algorithm progresses. Put differently, whenever any of these constraints are added to the FBBT at a given node, the current values for the bounds are used.
@@ 3.2.1 Numerical results
In order to assess the performance of the aforementioned approaches, we have studied a wide range of variations in how often FBBTOB and FBBTNB are applied along the branch-and-bound tree. We present here the best performing ones:
@@
FBBTOB: The OBBT-derived cuts from [Gleixner:2017] are added to FBBT at the root node and at nodes whose depth is a multiple of 50.
FBBTNB: The node by node cuts from [Ryoo:1996] are added to FBBT at the root node and at nodes whose depth is a multiple of 10.
FBBTOB+NB: Both approaches are applied together.
In Section A in the Appendix we report some additional results which, in particular, show that applying these enhancements at all nodes is not desirable. This is because the computational overhead required to perform FBBT’s constraint propagation on the extended set of constraints is not compensated by the gains from bound reduction. FBBTOB and FBBTNB provide a good balance between the computational effort and the reduction of the size of the branch-and-bound tree.
| (403 instances) | Solved (403) | Gap (111) | Time (134) | Pace (248) | Nodes (285) | Time(403) |
@@
| FBBTOB | 287 | |||||
| FBBTNB | 287 | |||||
| FBBTOB+NB | 287 |
@@ 3.3 Impact of the branching point on the branch-and-bound tree
Whenever a subproblem is solved at a node of the branch-and-bound tree, important decisions have to be made. Two crucial ones are the selection of the branching variable and of the precise value of that variable on which branching is done. As mentioned in the introduction, the choice of the branching variable has been thoroughly discussed by past literature, mainly for MILP problems (see, for instance, [Linderoth1999] and [Achterberg2005]), but also in the context of nonlinear optimization and spatial branching in [Tawarmalani2004] and [Belotti2009]. When it comes to polynomial optimization and the RLT technique in particular, [Gonzalez-Rodriguez:2023, Ghaddar:2023, Ghaddar:2025] offer deep analyses on the impact of variable selection.555In the seminal paper [Sherali:1992aa], starting from an optimal solution at a node in the branch-and-bound tree, the violation of each variable is computed as , where considers only the monomials present in PO. Then, the variable that maximizes this value is chosen for branching. More recent contributions rely on variations where the maxima in the computation of the values are replaced with weighted sums.
@@ 3.3.1 Numerical results
Given the results in Table 4, we see that OV and MP are particularly bad at Gap and Time, respectively. These results suggest that sticking to the “extreme” approaches, OV or MP, may be detrimental in performance with respect to the “interior” approaches. Put differently, would a hypothetical version of the solver capable of choosing the best performing approach between OV and MP outperform the strict convex combinations? The answer to this question is negative. Table 5 shows, for instance, that the aforementioned hypothetical configuration would be within 5% of the best performing approach in Time in, at most, 96 out of the 138 reported instances, whereas approach 0.5OV+0.5MP on its own is within 5% of the best performing one in 102 instances.
| (403 instances) | Gap5% (110) | Time5% (138) | Pace5% (250) | Nodes5% (280) |
@@ 3.4 Impact of combined enhancements
As a complement to the individual analysis in the preceding subsections, regarding the performance of the different approaches to enhance domain reduction techniques, we now study the impact of incorporating all of them together. Note that, although all the enhancements deal with different aspects of domain reduction, they are essentially independent from one another. Thus, a priori, there is no reason to expect that combining them might have a detrimental effect on performance: i) the enhancements in Section 3.1 only affect the OBBT at the root node, ii) the enhancements in Section 3.2 just affect the FBBT at all nodes, and iii) the enhancements in Section 3.3 deal with the branching point selection, which is independent of the OBBT and FBBT functionalities. For the sake of exposition, we just report the performance of combinations of the best performing approaches for each individual enhancement: SOCPM, FBBTOB+NB, and 0.5OV+0.5MP. Yet, it is important to note that retrieving dual information for (nonlinear) conic problems is far from straightforward. Thus, whenever a combination uses SOCPM, we replace FBBTOB+NB with FBBTNB, the best performing approach in Section 3.2 once we disregard the approaches applying FBBTOB.
@@
| SOCPM + FBBTNB | 287 | |||
| SOCPM + 0.5OV+0.5MP | 289 | |||
| FBBTOB+NB + 0.5OV+0.5MP | 287 | |||
| SOCPM + FBBTNB + 0.5OV+0.5MP | 287 |
In Table 6 we can see that, as expected from the previous results, the combinations that use approach 0.5OV+0.5MP are the most competitive ones. Among them, probably FBBTOB+NB + 0.5OV+0.5MP is slightly superior to the rest, with the raw 0.5OV+0.5MP performing very similarly. Interestingly, the combination SOCPM + 0.5OV+0.5MP is the one that solves most instances, despite not being so competitive in Time. It is worth noting that the approach that combines all enhancements, SOCPM + FBBTNB + 0.5OV+0.5MP, is not the most competitive one, suggesting that there may be some interdependence between the different enhancements after all.
@@ 4 Improving performance with machine learning
In this section we take the enhancements of Section 3 one step further. The goal is to study the potential for additional performance improvements by using machine learning techniques to predict the best combination of enhancements for a given instance. To this end, we also analyze the improvement on performance we would observe if we had an oracle signaling the best combination for each instance and, then, we use it as a reference to assess the quality of the learning.
@@ 4.1 Machine learning framework
In [Ghaddar:2023], the authors present a machine learning framework designed to “optimally” choose between different configurations of a solver when confronting a new instance, given its underlying features. In order to do so, the machine learning model is first trained on a predefined set of instances, using both their features and the performance of the different configurations on each of them. Table 7 presents the full list of the input variables (features).666VIG and CMIG stand for two graphs that can be associated to any given polynomial optimization problem: variables intersection graph and constraints-monomials intersection graph, and whose precise definitions is given in [Ghaddar:2023]. They capture diverse characteristics of each instance and are a key ingredient of the machine learning framework.
| Variables | No. of variables, variance of the density of the variables |
| Average/median/variance of the ranges of the variables | |
| Average/variance of the no. of appearances of each variable | |
| Pct. of variables not present in any monomial with degree greater than one | |
| Pct. of variables not present in any monomial with degree greater than two | |
| Constraints | No. of constraints, Pct. of equality/linear/quadratic constraints |
| Monomials | No. of monomials |
| Pct. of linear/quadratic monomials, Pct. of linear/quadratic RLT variables | |
| Average pct. of monomials in each constraint and in the objective function | |
| Coefficients | Average/variance of the coefficients |
| Other | Degree and density of PO |
| No. of variables divided by no. of constraints/degree | |
| No. of RLT variables/monomials divided by no. of constraints | |
| Graphs | Density, modularity, treewidth, and transitivity of VIG and CMIG |
For our analysis we follow [Ghaddar:2023]. We use Pace as the measure of performance (KPI) and use quantile regression forests for the learning [Meinshausen:2006], since in [Ghaddar:2023] they proved to be slightly superior to other techniques such as quantile generalized additive models and to stochastic gradient boosting for quantile regression. We want to emphasize that we have deliberately employed a simple learning approach, without any customization or adaptation to our specific setting. The fact that this approach yields significant improvements, particularly in Section 4.3, underscores the potential of such methodologies in this and related contexts. Yet, as we further discuss below, we acknowledge that it would be worth exploring how much farther these improvements might be pushed by using other learning methodologies such as deep learning and, in particular, graph neural networks [gupta2020], or enhancing the set of features with others that might be particularly relevant for domain reduction, e.g. enriching the features related to ranges of the variables and their interactions with other features.
One of the advantages of random forests, as well as other ensemble methods that use bagging, is that we the complete data set can be used to evaluate the model using the out-of-bag predictions, without randomly splitting the data into training and test sets.777As explained in [Ghaddar:2023], given a data set of size , “a large number of bootstrap samples of size are obtained from the data set by sampling with replacement. Each bootstrap sample leaves out, on average, about 37% of the observations (the left-out observations constitute the so-called out-of-bag sample). Then, individual trees are grown using these bootstrap samples. For each observation in the data set, the corresponding out-of-bag prediction is obtained by taking into account only those trees fitted on bootstrap samples that leave out that particular observation”. The learning is carried out on the same set of instances of the preceding section, using the statistical language R [R-Core-Team:2023] with the library ranger [Wright:2017].
@@
As mentioned above, not only we follow the approach in [Ghaddar:2023], but we also use the same set of features and the same parameters for quantile regression. More precisely, borrowing from [Ghaddar:2023], “learning is performed jointly on all sets of instances following the standard learning procedure. We randomly split the complete set of instances into two disjoint sets: the training set (70%) and the test set (30%). With the objective of obtaining a better performance, we gather the instances into “families” related to the groupings defined in the corresponding libraries, where those belonging to the same group share similar characteristics. The within-family proportion of instances is maintained through the splitting process. Moreover, in order to gain robustness, we construct 10 partitions of the dataset into training and test data and report aggregate results over all the partitions …when performing quantile regression one has to specify the -conditional quantile to be estimated.”888Given , let denote the estimated -conditional quantile of – a normalization of – for a given branching rule. This means that of the time the performance of that rule is expected to be larger than . In the numerical analysis we set , which proved to be the best performing value in [Ghaddar:2023], although the results were not particularly sensitive to this choice.
In the results’ tables in this section we report the performance of a special approach: Oracle, which represents the performance obtained when we choose, for each instance, the best configuration according to pace, i.e., it’s an ideal version representing what we would get if we had an oracle signaling the best approach for each instance. If, for a given instance, there are two options with the same pace, we choose the one with the smallest optimality gap.
@@ 4.2 Impact of machine learning on the individual enhancements
Table 8 summarizes the results of our learning methodology when applied individually to each one of the three different enhancements discussed in Section 3. More precisely, for each enhancement, the goal is to predict which one of the considered approaches will have a better performance on a given instance. To simplify the tables of results, among the original approaches we just represent the best performing one for the given enhancement, Best. “out-of-bag Q-RF” is the result of choosing, for each instance, the approach selected by the machine learning framework. Finally, “Improvement after learning” is the percentage of relative improvement from Best to out-of-bag Q-RF and “Oracle improvement” is the same but comparing Best to Oracle.
@@
| Oracle improvement | ||||
The results in Table 8 are somewhat mixed. For the conic bound tightening from Section 3.1, we see modest improvements in Gap and Time, but relatively far from the ones delivered by Oracle. Learning seems to perform particularly well for the FBBT enhancements from Section 3.2, with out-of-bag Q-RF significantly outperforming Best and getting around half-way of the improvements Oracle would achieve. Finally, regarding the selection of the branching point from Section 3.3, the learning does not seem to obtain much since, although it improves almost by 2% in Time with respect to Best, its performance deteriorates in Gap and Pace by 0.4% and 1.19%, respectively. It seems as if the inherent randomness behind the impact of the branching point on the generated branch-and-bound trees is hard to learn upon.
@@ 4.3 Impact of machine learning on the combined enhancements
Despite the mixed results of the previous section, one would expect that, when adding all of the enhancements and their combinations together, there would be more potential for learning, given the additional richness and diversity of the underlying approaches/configurations. For the sake of exposition, learning is not carried out on all approaches and combinations, but just on those reported in Section 3.4.
@@
The results in Table 9 show that machine learning performs fairly well. The learning approach outperforms Best (FBBTOB+NB + 0.5OV+0.5MP) according to all the metrics: it solves three more instances and obtains improvements of around 8% and 3% in Time and Gap. As expected, having a richer set of configurations to choose from also helps Oracle to obtain larger improvements.
@@
Figure 1 represents, side by side, how often each of the eight configurations is selected by out-of-bag Q-RF and by Oracle. We can see that out-of-bag Q-RF mimics quite well the behavior of Oracle. Importantly, Figure 1 also shows that a configuration with SOCPM is chosen in around one third of the instances, demonstrating the potential of using conic-based relaxations in the OBBT stages of global optimization solvers.
@@
Despite the apparently nice behavior of out-of-bag Q-RF displayed in Figure 1, it might even be the case that it never chose the best configuration and just happened to use the different configurations with the right frequencies. Figure 2 presents detailed information regarding the performance of the different configurations which, in particular, allows to assess how often each of them is the best one. The bar chart represents, for each configuration, the percentage of instances in which that configuration was the best one, the second best, and so on, for the eight ranking positions it might occupy depending on its performance relative to the rest.999Note that we have 8 configurations, in addition to out-of-bag Q-RF, and we just consider ranking positions from 1 to 8. This is because out-of-bag Q-RF always coincides with one of the eight configurations on which the learning is performed. Moreover, the numbers inside the bars indicate how close that configuration is, on average, to the best one for the instances in which it occupies the corresponding ranking position. More specifically, these numbers go from zero to one and are computed by dividing the best (smallest) Pace among the different configurations by the Pace of the current one, and then taking the average over the instances in each ranking position. Thus, Figure 2 allows to assess not only how often each approach is in each ranking position, but also how rapidly its performance deteriorates as it goes from the top positions to the bottom ones.
@@
Figure 2 shows that out-of-bag Q-RF comes on top more frequently than any other configuration, with 0.5OV+0.5MP close behind. Yet, the main difference between them comes from the fact that the relative performance of out-of-bag Q-RF deteriorates more slowly as it moves from being the best configuration to the second or third best, for example. In Figure 2 we can also see that the configurations that use SOCPM are the ones in which performance tends to deteriorate more quickly, which is consistent with the discussion in Section 3.1 about their “riskiness”. In this respect, since out-of-bag Q-RF uses SOCPM in nearly one third of the instances, the fact that it performs so well even when it does not choose the top performing configuration suggests that it is successful at learning to avoid using SOCPM in instances in which it does not perform well. Yet, the difference in performance between out-of-bag Q-RF and Oracle also suggests that there is room for further improvement. A natural direction would be to enhance the set of features of the instances, with the goal of capturing additional aspects that might be relevant for domain reduction.